# Exploring AI-based System Design for Pixel-level Protected Health Information Detection in Medical Images

### v3

@@ 4 Experiment
In this study, we assess various setups of a PHI detection pipeline by interchanging the roles of the presented models. Specifically, we study the four setups illustrated in Figure 2. The first setup employs three distinct models, each dedicated to one of the modules described in Section 3.1. This serves as a baseline setup, where each model performs the task it is specifically designed for. In the second setup, we assess the robustness of EasyOCR’s built-in object detection capabilities by using it for both text localization and extraction tasks. EasyOCR handles the detection of text regions and subsequently extracts the corresponding text. The third setup evaluates the language-vision capability of GPT-4o by assigning the text extraction step to it. In this configuration, GPT-4o receives cropped text regions, localized by the YOLOv11 model, performs OCR, analyzes, and classifies PHI content. The final setup explores GPT-4o as an end-to-end solution. GPT-4o takes full responsibility for all three pipeline steps: localizing text in the image, extracting the text, and analyzing it for PHI content. This setup, in particular, highlights a new use case where a single language model integrates vision and language tasks, demonstrating the potential for end-to-end applications.
For each of the introduced setups, we run the pipeline five times on the RadPHI-test dataset to assess the non-deterministic behavior of GPT-4o in generating responses. The final output of the PHI pipeline is evaluated in two stages. In the first stage, the pipeline is treated as a binary classifier for detecting PHI. The objective is to predict whether or not the image contains any PHI. In the second stage, only images identified as containing PHI are considered. The goal here is to evaluate whether the PHI categories are correctly classified. This stage functions as a multi-label classification task across eight PHI classes, see Table 1. In both stages, we calculate the mean values of accuracy, precision, recall, and F1 score by aggregating results from five distinct runs for each setup. The setup that performs best on the RadPHI-test dataset is then further evaluated on the MIDI dataset using the same evaluation metrics.
@@ 5.1 Rad-PHI dataset
(a) Precision
(b) Recall
(c) F1 Score
(d) Accuracy
Figure 7: Comparison of precision, recall, F1 score, and accuracy across four different setups, averaged from five separate runs of each setup obtained on the RadPHI-test dataset.
Precision Recall F1 Acc
Min Max Min Max Min Max Min Max
Global PHI 0.9875 0.9995 0.9482 0.9972 0.9701 0.9983 0.9506 0.9972
Date 0.9410 1.0000 0.8703 1.0000 0.9307 1.0000 0.9654 1.0000
Patient Name 0.6913 0.9908 0.9025 0.9969 0.7827 0.9938 0.8786 0.9968
Address 0.8725 1.0000 0.9571 0.9992 0.9128 0.9984 0.9563 0.9992
Gender 0.9160 0.9967 0.6934 0.9483 0.8178 0.9714 0.9245 0.9864
Email 0.9813 1.0000 0.8356 0.9964 0.9101 0.9985 0.9542 0.9990
Phone Nr 0.9648 0.9985 0.8339 0.9977 0.9056 0.9981 0.9548 0.9990
Age 0.9246 1.0000 0.6547 0.9913 0.7891 0.9957 0.9103 0.9978
Identifier 0.8026 1.0000 0.8886 0.9797 0.8685 0.9897 0.9359 0.9952
Table 3: Minimum and maximum scores of mean precision, mean recall, mean F1, and mean accuracy for each PHI category computed across all setups obtained on RadPHI-test. The color-coding indicates the setup attaining the corresponding metric value, i.e. Setup 1, Setup 2, Setup 3, Setup 4.
@@
1 YOLOv11 + EasyOCR + GPT-4o $0.080.08$ % 3,64136413,6413 , 641 s 1,686,27716862771,686,2771 , 686 , 277 227,702227702227,702227 , 702
2 EasyOCR + GPT-4o $0.100.10$ % 4,78547854,7854 , 785 s 1,704,06217040621,704,0621 , 704 , 062 259,819259819259,819259 , 819
3 YOLOv11 + GPT-4o $0.280.28$ % 12,2951229512,29512 , 295 s 5,639,64956396495,639,6495 , 639 , 649 425,935425935425,935425 , 935
4 GPT-4o $8.568.56$ % 3,43434343,$43$ , 434 s 3,249,45532494553,249,4553 , 249 , 455 250,577250577250,577250 , 577
@@
Table 4: Comparison of mean error rate, mean total time of analysis, the mean number of prompt and response tokens for 1,000 test cases across five runs for each setup evaluated on RadPHI-test. The pipeline error rate is the percentage of images that fail to process due to server issues or the presence of banned words in the requests sent to GPT-4o. The total time is estimated as the sum of the processing and analysis time required for the entire dataset.
@@
Figure 8: When text appears in multi-line blocks, out-of-the-box EasyOCR tries to fit bounding boxes that cover the entire text space. However, the bounding boxes might break the context consistency of the text. For example, Visit Date and 03-01-1989 are separated into two different crops, making it harder for the text analyzer to classify PHI content, particularly when a more fine-grain classification of dates is required, e.g., visit date or date of birth.
The results of the four setups are summarized in Figure 7, showing the precision, recall, F1 score, and accuracy in classifying images for global PHI (i.e., whether the image has PHI or not) and the corresponding eight PHI categories. Table 3 compares the pipeline error rate, total time taken to process 1000 images, and the number of prompt and response tokens used by GPT-4o across four setups. The pipeline error rate indicates the percentage of images that couldn’t complete the process due to GPT-4o’s inability to handle the request or the presence of its banned words in the query. Overall, all four setups demonstrate strong performance in deciding if an image contains PHI imprints, with all evaluation metrics related to the Global PHI label exceeding $0.940.94$. Setup 2 (EasyOCR + GPT-4o) and setup 4 (GPT-4o) exhibit slightly lower performance compared to the other setups. However, their performance remains robust, with the overall accuracy at $0.96040.9604$ and $0.95060.9506$, respectively. Setup 3 (YOLOv11 + GPT-4o) achieves the highest performance, with all four metrics surpassing $0.990.99$. This indicates that the combination of YOLOv11 for text localization and GPT-4o for text extraction and analysis excels at identifying images containing PHI. The overall high scores in all evaluation metrics, even in lower-performing setups, suggest that detecting the presence of PHI containing imprints is for all pipeline setups a solvable task.
Assessing the capability of different pipeline setups to perform fine-grained classification of PHI categories reveals strong performance deviations between the various setups. First, we observe that the PHI classes, such as date and address, are among the top-performing categories across all setups. This result is expected because both categories exhibit distinct semantic structures compared to other classes. On the other hand, gender information is the most challenging PHI class to detect, in particular when this information appears in abbreviated forms. Single-letter gender abbreviations often lack accompanying text like gender or sex, making detection more difficult. If the text extractor fails to recognize the exact character, the text analyzer is left without contextual clues to identify the PHI content and class. This issue is further complicated by potential confusion with single-character markers like L, R, A or P. Generally speaking, the bottleneck for fine-grained PHI category classification generally occurs during the text extraction step. When texts are incompletely extracted or when signaling texts are missed, it becomes significantly harder for the text analyzer to classify the content accurately. For example, Figure 8 illustrates how EasyOCR can break text continuity into separate bounding boxes, resulting in missing context and complicating the advanced analysis of PHI terms.
Among the four setups, Setup 2 has one of the lowest performance levels. This outcome is expected, as the detection component of EasyOCR is not trained on medical image data, making it less robust compared to a dedicated YOLOv11 model. Even when EasyOCR is used solely for text extraction (Setup 1), the final pipeline performance lags slightly behind Setup 3 where GPT-4o is used for integrated text extraction and analysis. Upon manual examination of specific cases, EasyOCR demonstrates more errors in text extraction compared to using GPT-4o for this task. However, in a majority of cases, GPT-4o, when used as a text analyzer, is powerful enough to infer the intended meaning from improperly extracted text by EasyOCR. This capability results in a comparable performance between Setup 1 and 3. While Setup 3 offers a slightly better overall performance, it has a significant drawback: the total analysis time is more three times that of Setup 1. This increase is primarily due to the latency associated with querying GPT-4o multiple times for extracting text from individual image crops. On average, the number of prompt and response tokens in Setup 3 is three times and two times higher than that in Setup 1, as shown in Table 4. Using GPT-4o for the entire pipeline is technically feasible, but it does carry a higher rate of pipeline errors and hallucinations. These errors frequently arise from unstructured outputs or the generation of special tokens that can cause the model to fail to respond. While our experimental result underscores the impressive vision capabilities of GPT-4o, it is still inadequate to entirely replace specialized models for text localization and extraction tasks in a single pass. For further evaluation on the MIDI dataset, we choose Setup 1 over Setup 3 as the gain in performance of the latter is very limited whereas inference costs, latency and token costs are much more favorable for Setup 1.
@@ 5.2 MIDI
In comparison to our simulated RadPHI-test dataset, the MIDI dataset is more challenging due to its increased diversity and a larger number of imprints. Relying on signaling words might not be sufficient to detect PHI effectively. For example, an image series by itself may not be classified PHI, but if it includes a date element, as shown in Figure 6, it is considered PHI. Overall, the performance of Setup 1 on MIDI is comparable to that obtained on RadPHI-test, see Figure 9 (left). Notably, the global PHI detection has a recall rate close to 1. Key categories demonstrating strong performance include dates, patient names, and addresses. However, similar to findings from the RadPHI-test, the identification of gender remains particularly challenging, especially in the context of abbreviations. The precision of the identifier category is diminished due to misclassifications of numerous image- and study-related identifiers as PHI. These misclassifications arise from our initial prompt to GPT-4o, which states that all forms of identifiers are regarded as PHI. To address the needs of this scenario, we refine the prompt by explicitly stating that study or image-related identifiers should not be classified as PHI. The improvement in precision for the identifier category following this prompt adjustment is demonstrated in Figures 9 (right) and  10. This case serves as a compelling illustration of how LLMs can adapt to detect PHI based on specific criteria, rather than being confined to a fixed definition. Overall, this allows users to configure the PHI detection system more flexibly, for example, to meet study-specific criteria.
(a) Same system prompt as in RadPHI-test
(b) Modified system prompt for identifier
Figure 9: Performance comparison of precision, recall, F1 score, and accuracy on the MIDI dataset using Setup 1: (a) performance with the RadPHI-test prompt and (b) performance with an adapted prompt for identifier classification.

### v4

@@ 4 Experiment
In this study, we assess various setups of a PHI detection pipeline by interchanging the roles of the presented models. Specifically, we study the four setups illustrated in Figure 2. The first setup employs three distinct models, each dedicated to one of the modules described in Section 3. In the second setup, we assess the robustness of EasyOCR’s built-in object detection capabilities by using it for both text localization and extraction tasks. EasyOCR handles the detection of text regions and subsequently extracts the corresponding text. The third setup evaluates the language-vision capability of GPT-4o by assigning the text extraction step to it. In this configuration, GPT-4o receives cropped text regions, localized by the YOLOv11 model, performs OCR, analyzes, and classifies PHI content. The final setup explores GPT-4o as an end-to-end solution. GPT-4o takes full responsibility for all three pipeline steps: localizing text in the image, extracting the text, and analyzing it for PHI content. This setup, in particular, highlights a new use case where a single language model integrates vision and language tasks, demonstrating the potential for end-to-end applications. For each of the introduced setups, we run the pipeline five times on the RadPHI-test dataset to assess the non-deterministic behavior of GPT-4o in generating responses. The final output of the PHI pipeline is evaluated at two levels: case and instance. At the case level, the focus is on whether the model can identify the presence of at least one PHI instance or a specific PHI class within the image. At the instance level, the evaluation checks whether the model accurately localizes every PHI instance or every instance of a specific class in the dataset. Performance is measured using recall, precision, false negatives (FN), and false positives (FP), which together provide insight into the model’s sensitivity in identifying PHI. A summary of the evaluation framework can be found in Table 3. In addition to the four pipeline variants, Microsoft’s open-source tool Presidio [16] is included in the benchmark. Presidio is designed for de-identifying images using NER models and regular expression patterns. We create two custom categories for patient names and identifiers using regular expressions, while also utilizing other built-in categories such as date, address, phone number, and email. Since Presdio is deterministic, we run it only once on our data and evaluate its performance. Our comparison with Presidio serves as a baseline for evaluating our LLM-based approaches. The setup that performs best on the RadPHI-test dataset is then further evaluated on the MIDI dataset using the same evaluation metrics.
Level Evaluation type Interpretation Metric Dataset
Case PHI presence Did the model correctly identify the presence of at least one PHI instance in the image? Recall, Precision, FN, FP RadPHI-test, MIDI
Case PHI class presence Did the model correctly identify the presence of at least one PHI instance of class X in the image? Recall, Precision, FN, FP RadPHI-test, MIDI
Instance PHI presence Did the model correctly localize every PHI instance in the dataset? Recall, Precision, FN, FP RadPHI-test(*), MIDI
Instance PHI class presence Did the model correctly localize every PHI instance of class X in the dataset? Recall, Precision, FN, FP RadPHI-test(*), MIDI
Table 3: Evaluation framework for PHI detection. The outputs are evaluated at two levels: Case-level, which considers whether the model detects the presence of at least one PHI instance or a specific PHI class in an image; and Instance-level, which evaluates the model’s ability to accurately localize each individual PHI instance or every instance of a specific PHI class across the dataset. (*) Instance-level evaluation is not performed in Setup 2 and Setup 4 because the coordinates are either not compatible to match with the ground-truth annotation (setup 2) or cannot be predicted (setup 4).
@@ 5.1 Rad-PHI dataset
(a) Case-level precision
(b) Case-level recall
(c) Instance-level Precision
(d) Instance-level Recall
Figure 7: Comparison of precision and recall at case and instance level. Metric averages from five separate runs of each setup on the RadPHI-test dataset. Setups 2 and 4 are not evaluated at the instance level because of incompatible or absent predicted imprint coordinates.
No. Images Mean Precision [FP] Mean Recall [FN]
with PHI Worst Setup Best Setup Presidio Worst Setup Best Setup Presidio
PHI Presence 779 0.9852 [11.4] 1.0000 [0] 0.9044 [54] 0.9017 [76.6] 0.9995 [0.4] 0.6560 [268]
Date 267 0.9309 [17.8] 1.0000 [0] 0.5734 [61] 0.8689 [35] 0.9963 [1] 0.3071 [185]
Patient Name 260 0.7532 [72.2] 1.0000 [0] 0.4102 [220] 0.8469 [39.8] 0.9923 [2] 0.5885 [107]
Address 248 0.9256 [17.8] 0.9959 [1] 0.7634 [31] 0.8855 [28.4] 1.0000 [0] 0.4032 [148]
Email 278 0.9950 [1.2] 1.0000 [0] 1.0000 [0] 0.8489 [42] 0.9849 [4.2] 0.2086 [220]
Phone Nr 260 0.9843 [3.2] 1.0000 [0] 0.9250 [6] 0.7631 [61.6] 0.9908 [2.4] 0.2846 [186]
Identifier 237 0.8301 [44.2] 1.0000 [0] 0.4065 [146] 0.8675 [31.4] 0.9949 [1.2] 0.4219 [137]
Table 4: Worst and best performing setup based on mean precision and recall at case level computed across five different runs on RadPHI-test (1,000 images). The mean False-Positive (FP) and False-Negative (FN) values indicated in the brackets are computed for the corresponding setup across five different runs as well. Note that Presido operates deterministically, with the scores reported derived from a single run. The color-coding indicates the setup attaining the corresponding metric value, i.e. Setup 1, Setup 2, Setup 3, Setup 4.
No. Instances Mean Precision [FP] Mean Recall [FN]
with PHI Worst Setup Best Setup Presidio Worst Setup Best Setup Presidio
PHI Presence 1550 0.9859 [21.6] 0.9869 [20.2] 0.3002 [1074] 0.9778 [34.4] 0.9808 [29.8] 0.3019 [1082]
Date 267 0.9963 [1] 1.0000 [0] 0.2985 [188] 0.9655 [9.2] 0.9963 [1] 0.2809 [192]
Patient Name 260 0.9642 [9.4] 0.9915 [2.2] 0.1511 [1117] 0.9731 [7] 0.9838 [4.2] 0.3923 [158]
Address 248 1.0000 [0] 1.0000 [0] 0.7619 [70] 0.9798 [5] 1.0000 [0] 0.3750 [155]
Email 278 1.0000 [0] 1.0000 [0] 0.9655 [2] 0.9496 [14] 0.9849 [4.2] 0.2014 [222]
Phone Nr 260 0.9961 [1] 1.0000 [0] 0.8235 [15] 0.9823 [4.6] 0.9908 [2.4] 0.2692 [190]
Identifier 237 0.9815 [4.4] 0.9913 [2] 0.2353 [312] 0.9671 [7.8] 0.9848 [3.6] 0.3038 [165]
Table 5: Worst and best performing setup based on mean precision and recall at instance level computed across five different runs on RadPHI-test (1,000 images). The mean False-Positive (FP) and False-Negative (FN) values indicated in the brackets are computed for the corresponding setup across five different runs as well. Note that Presido operates deterministically, with the scores reported derived from a single run. The color-coding indicates the setup attaining the corresponding metric value, i.e. Setup 1, Setup 3.
@@
1 YOLOv11 + EasyOCR + GPT-4o $0$ % 4,14341434,1434 , 143 s 1,727,29017272901,727,2901 , 727 , 290 280,998280998280,998280 , 998
2 EasyOCR + GPT-4o $0$ % 3,23532353,2353 , 235 s 1,744,12517441251,744,1251 , 744 , 125 298,087298087298,087298 , 087
3 YOLOv11 + GPT-4o $1.61.6$ % 7,68576857,6857 , 685 s 6,152,36961523696,152,3696 , 152 , 369 474,746474746474,746474 , 746
4 GPT-4o $5.185.18$ % 4,10441044,1044 , 104 s 3,530,30235303023,530,3023 , 530 , 302 287,229287229287,229287 , 229
Baseline Presidio $0$ % $288288$ s N/A N/A
@@
Table 6: Statistics related to pipeline execution: mean pipeline error (failure) rate, mean total time of analysis, the mean number of prompt and response tokens for 1,000 test cases across five runs for each setup evaluated on RadPHI-test. The pipeline error rate is the percentage of images that fail to process due to server issues or the presence of banned words in the requests sent to GPT-4o. The total time is estimated as the sum of the total processing and analysis time required for all 1,000 test cases.
@@
Figure 8: When text appears in multi-line blocks, out-of-the-box EasyOCR tries to fit bounding boxes that cover the entire text space. However, the bounding boxes might break the context consistency of the text. For example, Physician and Charlotte Lu are separated into two different crops, making it harder for the text analyzer to classify PHI content, particularly when a more fine-grain classification of names is required, e.g., physician or patient name.
The results of the four setups are summarized in Figure 7, showing the precision and recall in classifying images for PHI and the corresponding six PHI categories at the case and instance level. The worst and best setups based on precision and recall at two levels are summarized in Table 4-5.
PHI presence detection
Setups 1 and 3 demonstrate across all categories and metrics strong performance in detecting PHI presence at both the case and instance level. Setup 2, which relies on localization and text extraction by EasyOCR, exhibits a varying performance for different PHI categories. Setup 4, which utilizes GPT-4o end-to-end, obtains the lowest precision and recall. Setups 1 and 3 consistently achieve the highest results, attaining a maximum mean precision of 1.0 ± 0.0 (Setup 1) and a mean recall of 0.9995 ± 0.0007 (Setup 3) at the case level. Notably, Setup 3, with a case-level mean recall of 0.9995 (FN: 0.4 ± 0.55), indicates that, on average, fewer than one image containing at least one PHI is missed. However, with stricter evaluation at the instance level, the best precision and recall values decrease to 0.9869 ± 0.0008 and 0.9808 ± 0.0012 for Setup 1, resulting in an average of 29.8 ± 1.93 instances containing PHI being missed in the best-performing setup. Setups 2 and 4 are only assessed at the case level and show inferior performance compared to Setups 1 and 3. Additionally, Presidio falls behind all setups in both case and instance evaluations. At the case level, Presidio achieves precision and recall rates of 0.9044 and 0.6560, respectively, which drop significantly to 0.3002 and 0.3019 when evaluated at the instance level. This indicates that the tool has difficulty completely identifying all instances of PHI within the images.
PHI category detection
The performance of Setup 1 and Setup 3 is notably superior across all categories, while Setup 2 and Setup 4 significantly lag behind. Furthermore, Setups 1 and 3 demonstrate consistent performance across all metrics, in contrast to Setups 2 and 4, which exhibit considerable variability. Specifically, in terms of precision, Setups 2 and 4 achieve perfect scores for email and phone numbers but struggle with patient names and identifiers. Conversely, regarding recall, email and phone numbers experience the worst performance in Setup 2 and 4. Emails and phone numbers have unique formatting and signaling words, which allow easier differentiation from other types and hence higher precision. However, when texts are incompletely extracted or when key signaling elements are missed, it becomes challenging for the text analyzer to classify the content accurately. In these instances, email addresses and phone numbers can be easily confused with patient names and identifiers, as the former often include names while the latter consist of numbers. One of the reasons for incomplete text extraction is illustrated in Figure 8. EasyOCR in Setup 2 can break text continuity into separate bounding boxes, resulting in missing context and complicating the subsequent analysis of imprints. Similar to Setup 2 and Setup 4, Presidio also shows a high precision for selected PHI categories (phone numbers, email addresses) but low metric values on all other categories. In terms of recall, Presidio does not yield satisfactory results across all categories.
Tokens, latency, and pipeline error rate
Table 6 compares the pipeline error rate, total time taken to process 1,000 images, and the number of prompt and response tokens used by GPT-4o across four setups. The pipeline error rate indicates the percentage of images that couldn’t complete the process due to GPT-4o’s inability to handle the request or the presence of blacklisted words in the query. The blacklisted words are often encountered by erroneous OCR outputs, especially noisy or misrecognized names, can inadvertently resemble words that trigger these safety mechanisms. However, these encounters are rare compared to the general request processing error. While Setup 3 offers a slightly better overall performance in a few PHI categories, it has a significant drawback: the total analysis time is more than twice that of Setup 1. This increase is primarily due to the latency associated with querying GPT-4o multiple times for extracting text from individual image crops. On average, the number of prompt and response tokens in Setup 3 is three times and two times higher than that in Setup 1, as shown in Table 6. Using GPT-4o for the entire pipeline is technically feasible, but it does carry a higher rate of pipeline errors and hallucinations. These errors frequently arise from unstructured outputs or the generation of special tokens that can cause the model to fail to respond. In comparison to all four setups, Presidio has the shortest processing time (on average $0.2880.288$s per case), which is approximately a factor 10 faster than the LLM-based pipeline from Setup 2 (on average $3.2353.235$s per case). In regard to inference speed, lightweight or rule-based models, which do not require significant computational resources, clearly outperform our GPT-4o-based setup. It is important to note that our PHI pipeline has not yet been optimized for throughput. Smaller models and optimized inference processes will allow us to significantly reduce computation time in the future.
@@ 5.2 MIDI
For further evaluation on the MIDI dataset, we choose Setup 1 over Setup 3 as the gain in performance of the latter is very limited whereas inference costs, latency and token costs are much more favorable for Setup 1, see Table 6.
In contrast to our simulated RadPHI-test dataset, the MIDI dataset offers a realistic benchmark which is more challenging due to its increased diversity and a larger number of imprints. Relying on signaling words might not be sufficient to detect PHI effectively. For example, an image series by itself may not be classified PHI, but if it includes a date element, as shown in Figure 6, it is considered PHI.
When employing the same prompt used for RadPHI-test, we obtain the case-level precision and recall of 0.9509 ± 0.0127 and 0.9986 ± 0.0030, respectively as shown in Table 7 (a). On average, less than one image (FN: 0.2 ± 0.45) with PHI is overlooked. At the instance level where precise matching is crucial, precision and recall decline to 0.8797 ± 0.0096 and 0.9639 ± 0.0047, respectively. Key categories demonstrating strong performance include phone numbers, patient names, and dates. The precision of the identifier category is low due to misclassifications of numerous image- and study-related identifiers as PHI. The lower precision observed in the identifier category can be attributed to the misclassification of many image- and study-related identifiers as PHI. This issue stems from our initial guidelines for RadPHI-test, which regard all types of identifiers as PHI. To correct this, we refine the prompt by explicitly stating that study or image-related identifiers should not be categorized as PHI. The resulting enhancement in precision for the identifier category is reflected in Table 7 (b) and Figure 9. The effects of the new prompt are evident not only in the improvement across all categories at the instance level but also in a minor reduction in performance at the case level for categories such as patient name and phone number. While this observation is intriguing, we do not delve deeply into further prompt optimization at this stage. However, it warrants investigation in our future work, where we plan to evaluate additional LLMs on a larger dataset and study this phenomenon.
This prompt adaptation exemplifies how LLMs can be tailored to detect PHI in accordance with specific criteria, rather than adhering to a fixed definition. Ultimately, this flexibility allows users to configure the PHI detection system to better meet study-specific criteria.
No. Images / (a) Same prompt as for RadPHI-test (b) Modified prompt
Level Class Instances Precision [FP] Recall [FN] Precision [FP] Recall [FN]
Case PHI 147 0.9509 [7.6] 0.9986 [0.2] 0.9697 [4.6] (*) 1.0000 [0.0]
Date 116 0.9931 [0.8] 0.9914 [1.0] 1.0000 [0.0] (*) 0.9914 [1.0]
Patient name 104 0.9962 [0.4] 0.9981 [0.2] 1.0000 [0.0] 0.9769 [2.4] (*)
Address 36 0.8931 [4.0] 0.9222 [2.8] 0.9500 [1.8] (*) 0.9444 [2.0]
Phone Nr 21 0.9193 [1.8] 0.9714 [0.6] 1.0000 [0.0] (*) 0.9143 [1.8] (*)
Identifier 79 0.7427 [26.8] 0.9772 [1.8] 0.9398 [5.0] (*) 0.9848 [1.2]
Instance PHI 549 0.8797 [72.4] 0.9639 [19.8] 0.9721 [15.4] (*) 0.9781 [12] (*)
Date 207 0.9553 [8.8] 0.9082 [19] 0.9893 [2.2] (*) 0.9758 [5.0] (*)
Patient name 161 0.9851 [2.4] 0.9863 [2.2] 0.9776 [3.6] 0.9752 [4.9]
Address 39 0.8382 [7.0] 0.9282 [2.8] 0.9070 [3.8] (*) 0.9487 [2.0]
Phone Nr 24 0.9073 [2.4] 0.9750 [0.6] 0.9648 [0.8] (*) 0.9083 [2.2] (*)
Identifier 118 0.6589 [59.6] 0.9746 [3.0] 0.9325 [8.4] (*) 0.9814 [2.2]
Table 7: Performance comparison on the MIDI dataset using Setup 1: (a) performance with the RadPHI-test prompt and (b) performance with an adapted prompt for identifier classification. Superscripts (*) indicate statistical significance (pairwise Mann–Whitney U test, p < 0.05)